The revenue and cost equations for a product are
R=x(50-0.0002x) and C=12x+150,000
where R and C are measured in dollars and x represents the number of units sold. How many units must be sold be obtain a profit of at least $1,650,000?

Answer by stanbon(75887)

The revenue and cost equations for a product are
R=x(50-0.0002x) and C=12x+150,000
where R and C are measured in dollars and x represents the number of units sold. How many units must be sold be obtain a profit of at least $1,650,000?
---------------------
Profit = Revenue - Cost
P = x(50-0.0002x)-(12x+150,000)
----
P = 50x - 0.0002x^2 - 12x - 150000
-----
P(x) = -0.0002x^2 + 38x -150000
----
Solve: -0.0002x^2 + 38x -150000 >= 1650000
-------
etc.
===============
Cheers,
Stan H.
=================